OpenStudy (anonymous):

medals!!! The directions on a box of muffin mix calls for 1-2/3 cup muffin mix and 1/2 cups water. If there is only 1 cup of muffin mix left in the box, how much water should be used?

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

is there mult. choice?

OpenStudy (anonymous):

nope its just a fraction answer though

OpenStudy (superhelp101):

yep it is .8 or 4/5

OpenStudy (anonymous):

4/5

OpenStudy (superhelp101):

yea

OpenStudy (anonymous):

If 5 shirts cost $80, find the cost of 4 shirts.

OpenStudy (superhelp101):

set up a porportion

OpenStudy (anonymous):

x+5=80?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

????

OpenStudy (superhelp101):

\[\frac{ 5 shirts }{ 80 dollars} = \frac{ 4 shirts }{ x}\]

OpenStudy (superhelp101):

mult 80 and 4

OpenStudy (superhelp101):

then divide by 5

OpenStudy (anonymous):

320

OpenStudy (superhelp101):

320/5 is 64

OpenStudy (superhelp101):

so you were right good job!

OpenStudy (anonymous):

ok good got it right then